The perimeter of a square garden is 122 inches. How many centimeters are the length and width?

The perimeter of a square garden is 78ft. Find the length and width.

Answer by fcabanski(1391)

For a square P=4s where P is the perimeter and s is the length of a side.


Divide both sides by 4 to find that P/4 = s.


122/4=s=30.5 inches. The length and width of a square are the same. All the sides are the same length.


P=78=4s


Divide both sides by 4


78/4=s=19.5 ft. Every side of the square is 19.5ft. The length and width are both 19.5 feet.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com